No Work No Points Use Pencil Only Be Neat &Organized 1. Given P(A) = 0.65, P(B) = 0.55, and P(A and B) = 35, Find (a) (3 points) P(A or B) 1.2 (b) (3 points) P(AIB) o.Ss (c) (3 points) P(BIA)
Answer #1

SolutionA:

P(A or B)=P(A)+P(B)-P(A and B)

=0.65+0.55-0.35

=0.85

ANSWER:0.85

SolutionB:

P(A/B)=P(A and B)/P(A)=0.35/0.65=0.5384615

ANSWER:0.5384615

Solutionc:

P(B/A)=P(A/B)*P(B)/P(A)

=0.5384615*0.55/0.65

=0.45562

ANSWER:0.45562